OpenStudy (anonymous):

A bacteria culture begins with 13 bacteria which triple in amount at the end of every hour. How many bacteria are grown during the 10th hour? 13,312 26,624 255,879 2,302,911

OpenStudy (anonymous):

plzz help

OpenStudy (anonymous):

@ganeshie8

OpenStudy (dan815):

13*3 *3 *3....*3

OpenStudy (anonymous):

??

OpenStudy (dan815):

how many times did we triple till the 10th hour, you triple at the END of every hour

OpenStudy (anonymous):

9?

OpenStudy (dan815):

13*3 *3 *3....*3 = 13*3^9=?

OpenStudy (anonymous):

255,879

OpenStudy (anonymous):

tysm @dan815

OpenStudy (wolf1728):

That's worded pretty weird - when they say "how many are grown during the 10th hour" are they asking how many are present at the 11th hour?

OpenStudy (anonymous):

idek this is how it was on my quiz

OpenStudy (wolf1728):

So, when do you know if the answer is right or wrong?

OpenStudy (anonymous):

I submitted it and it was right:)

OpenStudy (wolf1728):

Okay so the answer is 255,879. Now we know.

OpenStudy (anonymous):

yep

OpenStudy (dan815):

yw @brebrebrebre